OpenStudy (anonymous):

let l>0 be a constant. Find the solution of the boundary value problem, where u"=0, u(0) = u(l) =0 ?

OpenStudy (anonymous):

\[u''=0\]integrate twice

OpenStudy (anonymous):

so, the result of it is Zero ?

OpenStudy (anonymous):

Im not quiet sure

OpenStudy (anonymous):

\[\int u'' dx=u'+c\]make sense?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

yes it does.. u reckon that what the question asked for ?

OpenStudy (anonymous):

yes \[u''=0\]\[\int u''dx=\int 0\]\[u'=c\]integrate both sides again\[\int u' dx=\int c \ dx\]\[u=cx+d\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

np :)